OpenStudy (anonymous):

Danny has a 15-foot board that he wants to use as a motorcycle ramp. If Danny wants the end of the ramp to be 4 feet off the ground, what angle will be formed between the board and the ground? Round the answer to the nearest whole degree.

OpenStudy (matt101):

Danny lifts one end of the board 4 feet off the ground. This creates a right triangle with a height of 4 feet and a hypotenuse of 15 feet. The angle between the board and the ground is given by this trigonometric equation: \[\sin \theta = \frac{4}{15}\]\[\theta = \sin^{-1}\frac{4}{15}\] Using a calculator I get about 15 degrees.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

@matt101